how would i verify the identity: (sect-1)/tsect= (1-cost)/t
Answer by stanbon(48564)

verify the identity: (sect-1)/tsect= (1-cost)/t
=====
Cross-multiply to get:
t(sec(t)-1) = tsec(t)(1-cos(t))
t(sec(t)) - t = tsec(t) - tsec(t)(cos(t))
-------------------------------------------
Subtract tsec(t) from both sides to get:
-t = -tsec(t)(cos(t))
----
-t = -t*1
-t = -t
=============================
Cheers,
Stan H.